OpenStudy (anonymous):

Chelsea left the White House and traveled toward the capital at an average speed of 34 km/h. Jasmine left at the same time and traveled in the opposite direction with an average speed of 65 km/h. Find the number of hours Jasmine needs to travel before they are 59.4 km apart.

OpenStudy (anonymous):

Gonna do one of these on your own at some point? ha

OpenStudy (anonymous):

dude i'm terrible at this

OpenStudy (anonymous):

i don't know why i don't get it

OpenStudy (anonymous):

Right tell me what variables we have?

OpenStudy (anonymous):

Just pluck out the numbers...

OpenStudy (anonymous):

chelseaspeed 34 km/h and jasmine 65km/h

OpenStudy (anonymous):

and the total distance is 59.4miles

OpenStudy (anonymous):

So we want the time for when the distance between Chelsea and Jasmine is 59.4km (not miles)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Ha, you got the answer for it?

OpenStudy (anonymous):

598.4=65(1.75+x)?

OpenStudy (anonymous):

thats where im at so far

OpenStudy (anonymous):

and no i havn't looked at the answer yet

OpenStudy (anonymous):

59.4**

OpenStudy (anonymous):

i meen instead of 598.4

OpenStudy (anonymous):

.84 hours?

OpenStudy (anonymous):

im not going to look at the answer till we figure this out

OpenStudy (anonymous):

Tell me what the answer is and I can see where I'm at haha. Not too sure of myself.. it's half 6 in the morning and I haven't slept ha.

OpenStudy (anonymous):

I get 36 minutes

OpenStudy (anonymous):

i got 3.2 hours

OpenStudy (anonymous):

ok h/o

OpenStudy (anonymous):

its .6 hours

OpenStudy (anonymous):

GIVE ME THE ANSWER :P

OpenStudy (anonymous):

0.6 hours is 36 minutes ;)

OpenStudy (anonymous):

haha nice

OpenStudy (anonymous):

My method is very complicated though... so take what you can from it. For every 1 km that Jasmine travels, Chelsea will travel 0.523 km. Therfore ratio of Chelsea to Jasmine is 523:1000. 59.4 /1523 = Distance travelled Chelsea distance travelled = 20.398 km ~ 20.4km Jasmine distance travelled = 39.002 km ~ 39.0km Time = Distance / Speed Jasmine Time = 39/65 Chelsea Time =20.4/34 Time for Chelsea and Jasmine should be the same, which if you work it out the calculations... They both equal 0.6 hours (36 minutes)

OpenStudy (anonymous):

im doing algebra 1 so i don't know it should just be some variation of d=rt